# PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have enhanced the efficacy of agent communication and social interactions. Despite these advancements, building LLM-based agents for reasoning in dynamic environments involving competition and collaboration remains challenging due to the limitations of informed graph-based search methods. We propose PLAYER*, a novel framework based on an anytime sampling-based planner, which utilises sensors and pruners to enable a purely question-driven searching framework for complex reasoning tasks. We also introduce a quantifiable evaluation method using multiple-choice questions and construct the WellPlay dataset with 1,482 QA pairs. Experiments demonstrate PLAYER*’s efficiency and performance enhancements compared to existing methods in complex, dynamic environments with quantifiable results.
@@ 1 Introduction
Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-as-Agent. Building upon this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023a), deception (Wang et al., 2023), and the spread of information (Park et al., 2023). Despite notable progress in enabling LLM-based agents to mimic human language, building agents for reasoning in dynamic environments, especially in scenarios considering competition and collaboration with other agents, remains a challenge, for example, the Murder Mystery Games (MMGs), a strategic game requires both cooperation and competition among 4-12 players through negotiation and tactical coordination (Figure 1). We argue that the primary reason is that most existing LLM-based agents involve the informed graph-based search method with a heuristics term to guide the reasoning progress. While this approach might be efficient in well-defined problems, where constructing a relation graph and producing a clear reasoning path as output is feasible, in complex problems requiring both cooperation and competition through natural language negotiation and tactical coordination, defining such a clear graph of potential actions becomes challenging. The main difficulty of applying traditional Multi-Agent Reinforcement Learning (MARL) methods in such scenarios arises from the challenges in defining state spaces, action spaces, and rewards, where the reward associated with the final decision often represents the only clear certainty (FAIR et al., 2022; Shi et al., 2024).
@@
Therefore, we propose a new framework based on an anytime sampling-based planner. Unlike informed graph-based search methods, such as A* (Dechter & Pearl, 1985), the anytime sampling-based planner, such as RRT* (Karaman & Frazzoli, 2011) or BIT* (Gammell et al., 2020a), tackles optimal motion planning problems in a more complex scenario. That is, by sampling possible states with detective sensors in real-time, the tree-style search path is constructed and pruned dynamically until the target is reached. Similarly, in MMGs, by proposing a set of sensors considering relations and emotions between players during interaction, along with a pruner that targets extracting highly suspicious murders, we construct a purely questioning-driven search framework. Our main contribution is that we proposed an approach which constructs a reasoning process without the need for pre-defined questions or prompt templates. Instead, it relies on a generator regulated by sensors and a pruner to efficiently generate high-quality questions for the agent’s reasoning task.
@@
Moreover, the current evaluation is difficult to quantify. These methods rely heavily on manual evaluation for in-depth understanding but are considerably influenced by subjective interpretation (Xu et al., 2023a; Wang et al., 2023).To address this, we propose a quantifiable and reproducible evaluation criterion by devising a series of multiple-choice questions. This approach encompasses three types of questions: fact-based questions that are automatically generated from an existing dataset focusing on character relationships (Zhao et al., 2024c), questions about the shared and individual objectives of each character set within the game, and questions probing the rationale behind the given responses.
@@
In summary, we made the following contributions: (1) We propose PLAYER*, a framework aimed at efficiently optimising path planning in MMGs. (2) We conduct experiments to assess the efficiency and performance enhancements brought by PLAYER*, demonstrating its superiority over existing multi-agent methods. (3) We propose a quantifiable evaluation method using multiple-choice questions focusing on facts, character objectives, and reasoning, mitigating the subjectivity in current evaluation practices with the construction and annotation of the corresponding dataset, WellPlay, which includes 1,482 QA pairs.  [^1] Our code and dataset are available at https://github.com/alickzhu/PLAYER.
Figure 1: The gameplay comprises three stages: introduction, discussion, and decision. During the introduction stage, players familiarise themselves with their characters’ background information and objectives, then briefly introduce themselves in character. Subsequently, they engage in interactive discussions to deduce the identity of the murderer.
@@ 2 PLAYER*
2.1 Problem Setting & Preliminary: Search by LLMs
In response to the complexities of social interactions such as MMGs, we have developed an innovative interactive framework tailored for such scenarios. As illustrated in Figure 1, this framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ for a game featuring $N_{a}$ playable characters. $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$ represents the set of victims in the game. Each agent $a_{i}$ is assigned to a corresponding character and initialised with its role background script $C_{i}$, suspicious state $s_{i}=[s_{ik}]_{k=1}^{N_{v}}$, and objectives $o_{i}$. Here, $C_{i}$ is designed from the unique viewpoint of that character, framing all relationships and events within the story from $a_{i}$’s perspective. The suspicious state $s_{ik}$ is a vector representing the suspicions of killing victim $v_{k}$ from the perspective of agent $a_{i}$, which is an $(N_{a}-1)$-dim vector, where the entries can be discrete or continuous based on suspicion strength. The complete game rules and examples can be found in the appendix A.
Unlike transitional planning task in which the state space can be accessed directly, $s_{i}$ in our task is generated from a language model. Therefore, searching in state space is done by adjusting the prompting. We define the possible prompts as an action space, allowing the states to be obtained through an LLM-based search: $s_{i}=P(C^{\prime}_{i},T)$, where $P$ is the LLM, $C^{\prime}_{i}\subseteq C_{i}$ represents the related scripts, and $T$ is the prompt instruction. The preliminary planner contains two main components:
  • Explorer: By choosing a different prompt template $T$, which is usually from a pre-defined prompt template set, explore the corresponding $s_{i}$.
  • Predictor: a self-reflection estimator to evaluate $s_{i}$. collaborate with other heuristic strategies (like A* or MC tree search) to determine if it is necessary to explore another $s_{i}$ by switching $T$, or using this $s_{i}$ and a clue to continuously explore in the next iteration.
In general, the preliminary method only construct a prompting based planner for a give script. However, in the complex scenario like MMGs, the objective can be multiple, noted as $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{i}}$, including both personal and shared objectives. Subsequently, the agents requires to engage in a simulated discussion by asking questions and expressing opinions. To ensure fairness, the game imposes a limit of $m$ questions per agent per round, spanning a total of $n$ rounds. Following these rounds, agents are required to make decisions and cast their votes to identify the suspected murderer. Therefore, we need to construct a more complicated strategies considering the simulation among agents.
@@ 2.2 Construction of Agents
The strength of the preliminary method is that the tree structure and the potential reasoning path is clearly defined by the $T$. However, the lack of flexibility and the over simplified explorer may lead to a over consistency result and limit the reasoning capability of LLM. But the entirely free generation, especially allowing the communication between agents may lead to an uncontrollable results. Therefore, in our framework, we propose another two components to regularise the generation of $T$ before and after prompting respectively:
  • Sensors: inspired by sociology, which claims that the interpersonal relationships are often conceptualised within a multidimensional space encompassing factors(Latané, 1981; Zhao et al., 2021; Trope & Liberman, 2010), we propose employing a set of sensors associated with each task. For example, the emotion, motivation, and suspicion, are considered as the hints to generate prompt for state searching.
  • Pruner: After prompting, we design a pruner to reduce the state space to only few high suspicions in the next round of reasoning.
Besides, To address the constraints of LLMs’ finite context window and the performance degradation associated with an increasing number of input tokens (Liu et al., 2023a), we implement a Memory Retrieval module to store and retrieve narrative scripts and dialog logs generated during gameplay. Specifically, we store the embeddings of all generated dialog logs and each agent’s script in a vector database dedicated to each agent. When an agent encounters a new event requiring action, we use the Faiss library (Douze et al., 2024) to retrieve relevant memories. This process enables the construction of prompts tailored to the predetermined maximum script length and dialog history length. It ensures consistent and coherent interactions among agents while assisting in their strategic planning.
Figure 2: PLAYER*’s planning strategy begins with an introduction round, during which each agent provides a brief self-introduction. Agents then initialise scenario understanding and begin questioning other agents. Subsequently, they start searching for further information by questioning other agents. Agents then adapt their level of suspicion and intended search area based on their received response, mirroring human gameplay strategies. This iterative process of inquiry and refinement continues until the final turn, when agents make decisions aligned with their objectives.
@@ 2.3 PLAYER* Planning Strategy
To navigate an unknown continuous space defined by natural language, which encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets, PLAYER* approximate the search domain through sampling, and planning the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions. As illustrated in Figure 2, this framework is fundamentally composed of two key components:
@@
Given the intricate social complexities inherent to MMGs, our proposed framework provides a robust solution through action generation with sensor components. Unlike prior approaches that required predefined prompts and a multitude of pre-set inquiries (Xu et al., 2023b), we generate actions directed towards the targets based on the values detected by those sensors. The task-specific sensors can vary across different tasks, and the only requisite input is a list of sensors. For MMGs, we have designated the input sensors $S$ to be $[S_{Emotion},S_{Motivation},S_{Suspicion}]$: Emotion, which assesses one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible; Motivation, evaluating from one’s perspective whether the individual possesses a motive to be the perpetrator; and Suspicion, objectively determining whether they had the opportunity to commit the crime.
@@
Since the number of inquiries an agent can pose is limited, each question incurs a cost, representing a missed opportunity for exploration, while the reward for such a cost is the information that can be obtained from the response. Under the assumption that increased inquiries to an agent corresponds to the heightened exploration within the world space related to that agent, the expected reward associated with questioning the same agent decreases in proportion to the number of inquiries posed. Therefore, it is necessary to evaluate the probability of obtaining valuable information by persisting in questioning a character who has already been asked many questions, even if this person might be highly suspicious. We estimate this probability $P_{informative}(D_{i},|Q_{i}|)$ using the previous dialog $D_{i}$ about the agent to be questioned, as well as the number of inquiries that have already posed $|Q_{i}|$. Specifically, for agent $a_{i}$ to be questioned, PLAYER* would generate the question $q$ using the values of sensors $S_{i}$ in conjunction with the expected reward $P_{informative}$ and other related information $(C^{\prime}_{i},D_{i})$, that is $f_{Q}:(C^{\prime}_{i},D_{i},S_{i})\rightarrow q$.
@@ 5                   $suspect\_list_{ik}$ = Suspect_Generation($s_{ik}$)
      // Constrain action space
@@
14                   $s_{ik}$ = Refinement_Action_Space($a_{i},a_{j},v_{i}$)
@@ 19       Result = Evaluate($a_{i}$)
Action Space Refinement with a Pruner
Following the search step, agents received the response from the questioned agent, as well as conversational exchanges among other agents. Each agent first updates its comprehension of the interpersonal relationships, thereby adjusting sensor values: $f_{sensor}:(C^{\prime}_{i},D_{i})\rightarrow S$. Subsequently, it also updated $P_{informative}$ based on the updated information $(S_{i},D_{i})$ and the changes in sensor values $\Delta S_{i}$.
In human gameplay, after an initial round of discussion, players often formulate judgments regarding the scenario and direct their focus towards a subset of highly suspicious individuals. This strategic refinement process is critical for conserving cognitive resources and enhancing the efficacy of inferential reasoning. Emulating the cognitive pattern of human beings, we introduce an Action Space Refinement strategy to refine the search space $f_{approximate}:(\mathcal{A},S_{i},P_{informative})\rightarrow s_{i}$: guided by sensors, the agent selects the most suspicious targets. We did not impose a hard restriction on the refinement process and allowed for flexible adjustment of the action space. For instance, the number of suspects could increase, or an agent might shift suspicion to someone previously overlooked. This is because new information might entirely overturn conclusions drawn in prior rounds, and we did not want to limit that possibility. This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads. By adopting this protocol, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and heightening the likelihood of identifying the perpetrator within the designated number of discussion rounds.

### v2

@@ Abstract
We propose PLAYER*, a novel framework that addresses the limitations of existing agent-based approaches built on Large Language Models (LLMs) in handling complex questions and understanding interpersonal relationships in dynamic environments. PLAYER* enhances path planning in Murder Mystery Games (MMGs) using an anytime sampling-based planner and a questioning-driven search framework. By equipping agents with a set of sensors, PLAYER* eliminates the need for pre-defined questions and enables agents to navigate complex social interactions. We additionally make a contribution by introducing a quantifiable evaluation method using multiple-choice questions and present WellPlay, a dataset containing 1,482 question-answer pairs. Experimental results demonstrate PLAYER*’s superiority over existing multi-agent methods, enhancing the generalisability and adaptability of agents in MMGs and paving the way for more effective multi-agent interactions.
PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games
Qinglin Zhu1∗, Runcong Zhao1∗, Jinhua Du2, Lin Gui1, Yulan He1,3 1King’s College London, 2Huawei London Research Centre, 3The Alan Turing Institute {qinglin.1.zhu, runcong.zhao}@kcl.ac.uk{jinhua.du}@huawei.com {lin.1.gui,yulan.he}@kcl.ac.uk
@@ 1 Introduction
Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-based agents (Soni et al., 2023; Cherakara et al., 2023). Building on this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023), deception (Wang et al., 2023), and the spread of information (Park et al., 2023). Despite these advances, building agents that intentionally interact with humans in dynamic environments remains challenging due to several limitations. First, current state-of-the-art approaches struggle with complex questions that lack straightforward answers and have difficulty in understanding interpersonal relationships. As a result, in addition to the game rules, agents often require additional human-written guides tailored to specific scenarios, relying on human context understanding (Xu et al., 2023; Wang et al., 2023). This limitation highlights the need for significant manual efforts and reduces the generalisability of these approaches across different domains and situations. Second, applying traditional Multi-Agent Reinforcement Learning (MARL) methods in such scenarios is challenging because of the difficulties in defining state spaces, action spaces, and rewards. The reward associated with the final decision (e.g., identifying the murderer) often represents the only clear certainty (win or lose) (FAIR et al., 2022; Shi et al., 2024). Third, current evaluation methods are difficult to quantify. These methods rely heavily on manual evaluation to gain a deep understanding, but they are considerably influenced by subjectivity (Xu et al., 2023; Wang et al., 2023).
Figure 1: The murder mystery game unfolds as players take on roles, question suspects based on their character’s objectives, and use deduction to identify the killer, while achieving their ideal role-specific ending. The complete game rules can be found in the appendix A.
To address the limitation of lacking generalisability, we propose a novel framework that efficiently optimises path planning in Murder Mystery Games (MMGs), a strategic game involving both cooperation and competition among 4-12 players through natural language negotiation and tactical coordination (as illustrated in Figure 1). Our approach eliminates the need for pre-defined questions or instructions, allowing agents to play various scripts without extra human-written guides, just like human players. This framework enables agents to navigate complex social interactions and make decisions by considering the types of sensors they need. In the case of MMGs, this involves a set of sensors that detect emotions between players, motivations, and opportunities for actions like murder.
@@
To address the difficulty in defining state spaces and rewards, we designed an anytime sampling-based planner instead of using MARL. Unlike informed graph-based search methods, such as A* (Dechter and Pearl, 1985), our planner, inspired by algorithms like RRT* (Karaman and Frazzoli, 2011) and BIT* (Gammell et al., 2020), samples possible states with detective sensors in real-time. The tree-style search path is constructed and pruned dynamically until the target is reached. For MMGs, we combine sensors with a pruner that focuses on extracting highly suspicious murders to construct a purely questioning-driven search framework. PLAYER* is guided by sensors to efficiently generate high-quality questions to uncover more clues, and uses a pruner to more precisely identify the killer.
@@
To address the evaluation problem, we propose more quantifiable and reproducible evaluation criteria by devising a series of multiple-choice questions. This approach encompasses three types of questions: (1) Objective questions about the shared and individual objectives of each character in the game, (2) Reasoning questions probing the reasoning behind the given responses, and (3) Relations questions that are automatically generated from an existing dataset focusing on character relationships (Zhao et al., 2024c). In summary, we have made the following contributions:
@@
(1) We propose PLAYER*, a framework that enhances the generalisability and adaptability of agents by efficiently optimising path planning in MMGs. (2) We designed an anytime sampling-based planner to dynamically construct and prune the search path, enhancing performance and efficiency. (3) We introduce a quantifiable evaluation method using multiple-choice questions focusing on character objectives, reasoning, and facts, thereby reducing the subjectivity in current evaluation practices. We also constructed and annotated a corresponding dataset, WellPlay, which includes 1,482 QA pairs. [^1] Our code and dataset are available at https://github.com/alickzhu/PLAYER.
@@ 2 PLAYER*
In this section, we describe the MMG settings and provide an overview of how PLAYER* works.
2.1 Problem Setting
Figure 2: Problem Setting.
Figure 3: PLAYER* begins with an introduction round, where each agent provides a brief self-introduction. Following this, agents initialise scenario understanding and begin to search for further information by questioning other agents. We use arrows to identify their questioning actions, with blue arrows representing questions asked by the agent itself and yellow arrows representing questions asked by other agents, which are also remembered by the given agent. Based on the responses they receive, agents update their suspicion state, identified by the black arrow (which decides who to ask), and the corresponding action space, identified by the pink area (which decides what questions to ask). The iterative process of inquiry and refinement continues until the final turn, when agents make decisions aligned with their objectives.
In response to the complexities of social interactions in settings such as MMGs, we have developed an innovative interactive framework tailored for such scenarios. As illustrated in Figure 2, this framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ for a game with $N_{a}$ playable characters and a set of victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$. Each agent $a_{i}$ is assigned to a character and initialised with its role background script $C_{i}$, suspicious state $s_{i}=[s_{ik}]_{k=1}^{N_{v}}$, and objectives $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{i}}$. The script $C_{i}$ is crafted from the unique viewpoint of that character, framing all relationships and events within the story from $a_{i}$’s perspective. The suspicious state $s_{ik}$ is represented as an $(N_{a}-1)$-dimensional vector that encodes agent $a_{i}$’s beliefs about other players’ suspicions regarding the murder of victim $v_{k}$. The vector entries can be either discrete ($\in{\{0,1}\}$) or continuous values ($\in[0,1]$), reflecting the strength of the suspicions. Unlike traditional planning tasks in which the state space can be accessed directly, $s_{i}$ in our scenario is generated from a language model. We define the possible questions as an action space, allowing the derivation of suspicious states through an LLM-based search. For an illustrative example, refer to Appendix A.3.
As shown in Figure 3, each agent is provided with character-specific information, enabling them to interact with others based on their scripts and historical interactions within the game. Subsequently, agents engage in simulated discussions by asking questions and expressing opinions. To ensure fairness, each agent is limited to asking $m$ questions per round over $n$ rounds. After the discussion, agents must collectively decide and vote on the suspected killer.
@@ 2.2 Construction of Agents
To enable agents to think and act like human players, we equip them with the following components: (1) Game Rules. The essential information for both human and AI players. (2) Chain-of-Thought. Facilitating a sequential reasoning process to aid in problem-solving. (3) Sensors. In sociology, interpersonal relationships are often conceptualised within a multidimensional space encompassing factors like time, distance, and various other elements (Latané, 1981; Zhao et al., 2021; Trope and Liberman, 2010). Drawing inspiration from this concept, we propose employing a set of sensors associated with each task. For example, in a murder mystery game, these factors could include emotion, motivation, and suspicion. (4) Memory Retrieval. To address the constraints of LLMs’ finite context window and the performance degradation associated with an increasing number of input tokens (Liu et al., 2023a), we implement this module to store and retrieve narrative scripts and dialog logs generated during gameplay. Specifically, we store the embeddings of all generated dialog logs and each agent’s script in a vector database dedicated to each agent. When an agent encounters a new event requiring action, we use the Faiss library (Douze et al., 2024) to retrieve relevant memories. This process enables the construction of prompts tailored to the predetermined maximum script length and dialog history length. It ensures consistent and coherent interactions among agents while assisting in their strategic planning throughout the game. (5) Self-Reflection. Preserving historical experience by reflecting on the current situation to provide a more detailed, accurate, and profound answer.
@@ 2.3 PLAYER* Planning Strategy
To navigate an unknown continuous space defined by natural language, which encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets, PLAYER* approximates the search domain through sampling, and plans the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions. As illustrated in Figure 4, this framework is fundamentally composed of two key components:
Figure 4: Search and Approximate. PLAYER* generates questions based on sensor-detected values and identifies the killer more precisely with a prunner. This dynamic question generation allows agents to adjust to the specific circumstances of each game session, ensuring a more focused and effective investigation.
@@
Given the intricate social complexities inherent to MMGs, our proposed framework provides a robust solution through action generation with sensor components. Unlike prior approaches that required predefined prompts and a multitude of pre-set inquiries (Xu et al., 2023), we generate actions directed towards the targets based on the values detected by those sensors. The task-specific sensors can vary across different tasks, and the only requisite input is a list of sensors. For MMGs, we have designated the input sensors $S$ to be $[S_{Emotion},S_{Motivation},S_{Opportunity}]$: Emotion: This sensor assesses an agent’s disposition towards another character, reflecting one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible. Motivation: This sensor evaluates from one’s perspective whether an individual possesses a motive to be the perpetrator. Opportunity Assessment: This sensor objectively determinies whether a character had the opportunity to commit the crime.
The sensor values are obtained based on the related information retrieved from the script and dialog, represented by the function $f_{sensor}:(C^{\prime}_{i},D_{i})\rightarrow S$, where $C^{\prime}_{i}\subseteq C_{i}$ represents the related scripts and $D_{i}$ denotes the associated dialogue history. Then, for agent $a_{i}$ to be questioned, PLAYER* uses LLMs to sample the question $q$ from the action space given the values of sensors $S_{i}$ and related information $(C^{\prime}_{i},D_{i})$, represented by the function $f_{Q}:(C^{\prime}_{i},D_{i},S_{i})\rightarrow q$.
Action Space Refinement with a Pruner
@@
Following the search step, agents received the response from the questioned agent, as well as conversational exchanges among other agents. Each agent first updates its understanding of interpersonal relationships, thereby adjusting sensor values $S$. Subsequently, it also updated $P_{informative}$ based on the revised information and changes in sensor values. Since the number of inquiries an agent can pose is limited, each question incurs a cost, representing a missed opportunity for exploration. The reward for questioning is the information that can be obtained from the anticipated response. As the number of inquiries directed at an agent increases, the assumed exploration within the world space relevant to that agent also rises. Consequently, the expected reward associated with questioning the same agent diminishes proportionally to the number of inquiries already made. Therefore, it becomes essential to evaluate the probability of obtaining valuable information by persisting in questioning a character who has already been asked many questions, even if they appear highly suspicious. We estimate this probability $P_{informative}(D_{i},|Q_{i}|)$ using the previous dialog $D_{i}$ concerning the agent to be questioned, and the number of inquiries already posed $|Q_{i}|$.
In human gameplay, after an initial round of discussion, players often formulate judgments regarding the scenario and direct their focus towards a subset of highly suspicious individuals. This strategic refinement process is critical for conserving cognitive resources and enhancing the efficacy of inferential reasoning. Emulating the cognitive pattern of human beings, we refine the search space $f_{approximate}:(\mathcal{A},S_{i},P_{informative})\rightarrow s_{i}$: guided by sensors, the agent selects the most suspicious targets (See Lower Part of Figure 4).
We did not impose a hard restriction on the refinement process and allow for flexible adjustment of the suspicious state. For instance, the number of suspects could increase, or an agent might shift suspicion to someone previously overlooked. This is because new information might entirely overturn conclusions drawn in prior rounds, and we did not want to limit that possibility. Then we update the action space based on the suspicious state. This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads. By adopting this protocol, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and heightening the likelihood of identifying the perpetrator within the designated number of discussion rounds.
@@ 5                   $suspect\_list_{ik}$ = Suspect_Generation($s_{ik}$)
      // Update the suspect list
@@
14                   $s_{ik}$ = Update_Suspicious_State($a_{i},a_{j},v_{i}$)
@@ 19       Result = Evaluate($a_{i}$)
3 The WellPlay Dataset
Script Agents Victims #token(CN) #token(EN) Question
avg overall avg overall Objective Reasoning Relations overall
Death Wears White 9 1 3,191 28,716 1,742 15,681 10 102 72 184
Ghost Revenge 7 3 5,488 38,415 3,960 27,723 19 152 69 240
Danshui Villa 7 2 5,111 35,779 3,339 23,370 12 128 63 203
Unfinished Love 7 2 2,501 17,507 1,652 11,562 12 61 72 145
Cruise Incident 5 1 1,263 6,313 808 4,040 4 24 30 58
Sin 4 1 2,121 8,485 1,378 5,512 3 20 21 44
Deadly Fountain 4 1 1,852 7,410 1,194 4,775 3 21 12 36
Unbelievable Incident 5 1 3,182 15,912 2,012 10,062 4 24 15 43
Desperate Sunshine 4 1 3,370 13,481 2,219 8,874 3 18 36 57
Riverside Inn 4 1 1,910 7,638 1,257 5,028 3 18 18 39
Solitary Boat Firefly 6 4 8,894 53,362 6,874 41,244 20 109 69 198
Manna 6 3 9,028 54,169 6,492 38,954 24 123 88 235
Avg 5.67 1.75 3,993 23,932 2,744 16,402 9.75 66.67 47.08 125.50
Sum 68 21 47,911 287,187 32,927 196,825 117 800 565 1482
Table 1: Dataset Statistics. Agents is the count of players, Victims is the number of script victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively. Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types. The number of evaluation questions varies based on script complexity, with more complex scripts generating a larger volume of questions.
WellPlay Dataset
We built an evaluation dataset, called WellPlay, from the existing Conan dataset (Zhao et al., 2024c), derived from background narratives created for MMGs. Conan labels relationships from each character’s perspective, along with ground truth relationships, consolidating all available information. To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information to reduce controversial answers. Besides existing labels of interpersonal relationships that gauge agents’ understanding of the situation, we also recruited four annotators to label character objectives and reasoning. The dataset encompasses 12 MMGs, comprising a total of 1,482 evaluation questions. On average, each game features 5.67 agents and 1.75 victims as detailed in Table 1 . Our evaluation questions are grounded in characters’ objectives in the game and are crafted with input from human players, focusing on the following question types:
(1) Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles wallet, for each character in the game.
(2) Reasoning. This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, disappearance, explosion); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, financial conflicts, manslaughter).
(3) Relations. This includes interpersonal relationships between victims and other characters, as well as relationships among suspects.
Evaluation Metrics
We report the performance of each type of question and also calculate the agent’s final score as Awarded PointsTotal $\frac{\text{Awarded Points}}{\text{Total Points}}$ to determine its overall performance. The awarded points follow the scoring system used in MMGs, which awards 10 points for objective questions, 5 points for reasoning questions, and 2 points for relation questions. For these calculations, we utilise weighted mean and weighted standard deviation methodologies. In addition to assessing the accuracy of the three types of questions and the overall score, we also evaluate the Win Rates of players for performance assessment. The game rules are as follows: every player, including the murderer, can vote, and the player with the majority number of votes ($\geq 50\%$) will be eliminated. Additionally, we assume that murderers would always vote for someone else instead of voting for themselves. If the true murderer is successfully voted out, the game is considered a victory for the players seeking to identify the culprit. The detailed about the dataset and calculations is provided in Appendix B.